OpenStudy (anonymous):

s=(n-2)180 solve for n

OpenStudy (anonymous):

distribute the 180: 180n - 360 = s Add 360 to both sides: 180n = s + 360 Divide both sides by 180: n = (s+360)/180

OpenStudy (anonymous):

thank you. I thought it was n=s+2 but based on this you are saying the s+360 are over the 180 and you cannot break that down further?

OpenStudy (anonymous):

The equation could be written n = (s/180) + (360/180) n = (s/180) + 2 but NOT (n =s + 2)<---bad

OpenStudy (anonymous):

thank for your help

OpenStudy (anonymous):

No problem if you get confused you can always try out numbers: Say s = 2 does (2 + 360)/180, 362/180 equal 2 + 2? no way So it is very easy to check if your confused